(1 point) In an experiment, a ball is drawn from an urn containing 10 red balls and 14 purple balls. If the ball is red, three coins are tossed. Otherwise two coins are tossed.
How many elements of the sample space will have a red ball?

How many elements of the sample space are there altogether?
Answer #1

a) Number of possible outcomes with tossing 3 coins = 2*2*2 = 8

Hence,

Number of elements in sample space that will have a red ball = 8

b) Total elements in the sample space

= Elements with red ball + Elements with Purple ball

= 2*2*2 + 2*2

= 8 + 4

= 12